# Schrödinger Bridge Matching for Tree-Structured Costs and Entropic Wasserstein Barycentres

### v1

@@ D.4 Higher-dim Gaussian Experiments
For TreeDSBM, we use $\varepsilon=0.0001$ and run for 4 IMF iterations. For training the vector fields, we use 10,000 training steps and a batch size of 4096. We use learning rate 1e-3 and exponential moving average parameter of 0.99. We generate a batch of 50,000 training couplings for subsequent TreeIMF iterations (10,000 simulated from each marginal).
@@
D.5 Ave! Celeba Benchmark
We use the Ave! Celeba data provided in Korotin et al. [2022], which consists of around 67k samples in each marginal, and each image is shape $64\times 64\times 3$.
It is known that performing bridge-matching between complex datasets such as images can be challenging, so for the first step we instead pretrain models using single bridge-matching iteration from a standard Gaussian to each marginal. To obtain the next coupling for training, we run the process from one of of the marginals to the latent representation in the Gaussian, and then out to the other marginals. This aids in learning, as there is often good structure preserved between the obtained samples from each marginal. For subsequent iterations we also warmstart the parameters from these pretrained models. The experiments were conducted on Nvidia A100 GPUs on Google Colaboratory.
Hyperparameters
We use the UNet architecture of Song et al. [2023], with 128 channels, channel multiples of (1,2,2,2), attention at layers (32,16,8), and 4 residual blocks at each layer.
We use $\sigma=0.01$ and train each bridge-matching procedure with batch size 32, at a learning rate of 1e-4 and with exponential moving average weight of 0.9997. For pretraining, we run for 100,000 training steps (which takes approximately 12 hours), and for subsequent IMF iterations we run for 20,000 steps (which each take around 2.5 hours).
Figure 8: Additional samples from TreeDSBM applied to the Ave! Celeba benchmark. Many samples are transported well, but some pick up unwanted artifacts. We discuss the findings from this experiment below.
Discussion of Results
TreeDSBM is able to scale to the high-dimensional setting well, but the final samples do not match the performance of state-of-the-art results, for example in Kolesov et al. [2024a]. Observe that some of the generated samples are good, but some contain additional artifacts that should not be present. This is due to the initial pretraining coupling; when generating samples $Y_{i}$ from the pretraining coupling, we simulate from one marginal to the Gaussian latent, and then our to the other marginals. This often results in strong structural similarities between the obtained $Y_{i}$ which give good barycentre samples for training the next IMF iteration. However, sometimes the coupling samples $Y_{i}$ do not resemble each other, and the resulting barycentre sample consists of separate overlaid images. This can be difficult for the algorithm to recover from, and results in the artifacts visible in some of the generated images.
Overall this suggests a limitation of TreeDSBM for this particular benchmark—the true transport maps are in fact very simple in this specific example (primarily just colour changes), but it is difficult for TreeDSBM to learn this because communication between the edges is infrequent and only occurs after each IMF iteration. In contrast, state-of-the-art methods for this benchmark optimise all the maps together and with much more interaction between them, which we anticipate is a better inductive bias for the shared structure present in this specific problem setting. Note that the fact that TreeDSBM optimises the edges separately is in fact a strength of the approach in many settings; it results in stable training without needing adversarial objectives, and allows for speed-ups by training the edges simultaneously.
We also anticipate that the performance of TreeDSBM in this setting could be improved through architectural changes (to promote suitable inductive biases for this specific problem) and other implementation tricks from the flow-matching literature (for example, using preconditioned flow parameterisations [Karras et al., 2022], and techniques to mitigate marginal drift in reflow methods [Kim et al., 2025]). Such investigations offer promising directions for future work.
@@ D.5 Ave! Celeba Benchmark
[…]
@@ Appendix E Additional Experiments
E.1 Beyond Star-Shaped Trees
@@
So far, we have only demonstrated the empirical performance of TreeDSBM on star-shaped trees, as we have focused on computing Wasserstein barycentres. Finally, we demonstrate that our TreeDSBM also works for non-star shaped trees, and thus has potential applications beyond only barycentre computation. We consider a simple 2-dimensional example with the same tree structure as shown in the TreeIMF diagram in Figure 1, with standard scikit-learn distributions on the observed leaves and each edge having length 1. As in the previous 2-dimensional barycentre problem, this is a challenging task due to the discontinuous transport maps and fragmented supports at the solution. We run TreeDSBM for 4 IMF iterations with $\varepsilon=0.1$, and train each edge for 10,000 iterations with learning rate 1e-3 and exponential moving average parameter of 0.99. We plot the obtained measures in Figure 9, and see that TreeDSBM is again able to learn the complex mappings required for this setting. While direct applications of non-star-shaped trees are less clear than in the barycentre case, examples have been studied in Haasler et al. [2021], Solomon et al. [2015], and they could have potential applications for modelling temporal behaviour of population dynamics. We leave investigating possible applications of general tree-structured costs for future work.

### v2

@@ D.4 Higher-dimensional Gaussian experiments
    For TreeDSBM, we use $\varepsilon=$1e-4 and run for 4 IMF iterations. For training the vector fields, we use 10,000 training steps and a batch size of 4096. We use learning rate 1e-3 and exponential moving average parameter of 0.99. We generate a batch of 50,000 training couplings for subsequent TreeIMF iterations (simulated equally from each marginal). Results reported for TreeDSBM are the average over 5 runs.
@@
    Convergence speed
    We provide the values of the $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ metric as the IMF iterations progress in Table 8, and again we observe very fast convergence.
[…]
    IMF 1 IMF 2 IMF 3 IMF 4
    $\downarrow\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ 16.0 $\pm$ 0.01 0.12 $\pm$ 0.01 0.13 $\pm$ 0.02 0.14 $\pm$ 0.03
    $\downarrow L^{2}\mathrm{-UVP}$ 16.7 $\pm$ 0.01 1.19 $\pm$ 0.02 1.18 $\pm$ 0.02 1.18 $\pm$ 0.03
@@
    The best choice of barycentre algorithm may depend on the specific problem setting. For example, if the transport maps exhibit complex behaviour (possibly due to lower-dimensional, manifold-like structures in the datasets) then the flow-based approach of TreeDSBM will likely perform strongly (such as in the 2dd barycentre experiment in Section 4). Also, when fast training is required then our experiments suggest TreeDSBM is a strong option. On the other hand, if fast inference is important then a one-step-generation solver such as NOTWB might be preferable. Note that one could incorporate distillation techniques from the flow-matching literature for improving the speed of TreeDSBM inference after training.
@@ Appendix E Additional experiments
    E.1 Further comments regarding computational considerations
    Choice of entropy regularisation
    Choosing the entropy regularisation parameter $\varepsilon$ is a perennial question in entropic OT. Standard methods to choose this value in commonly used OT libraries (for example, choosing in proportion to the costs) provide good guidance for choosing suitable values. Typically, one may want to choose $\varepsilon$ as small as possible the reduce the entropic bias. One advantage of TreeDSBM over TreeDSB is that it allows for much smaller epsilon (TreeDSB does not converge for too-small $\varepsilon$, as simulated trajectories struggle reach the other marginals). We provide a visualisation of the role of $\varepsilon$ in the 2dd example in Figure 6, and also for two values of $\varepsilon$ for the MNIST experiment in Figure 7. In Table 10, we also add some further quantitative results for different $\varepsilon$ values, in the 2dd and subset posterior aggregation settings.
    Fitting to the marginals
    One of the limitations of our approach is that errors can accumulate in the marginals as IMF iterations proceed; this is a limitation inherited from standard IMF and similar reflow methods. Standard techniques from the literature can be used to mitigate this (such as rotating the starting marginal as in [77], or using the projection methods in [42]). It is therefore important that the bridge-matching steps fit the marginals accurately, and hyperparameters should be chosen accordingly. To provide an indication of how the learned bridge-matching quality affects the overall solution, we provide results for varying batch size on the 2dd and data aggregation experiments in Table 11, for the hyperparameters used in the paper. We have also provided results assessing how the fit to the marginals changes as the IMF iterations progress in the 2dd and MNIST experiments in Tables 6 and 7 respectively.
    Table 10: Effect of entropy-regularisation parameter $\varepsilon$ in the 2dd and data aggregation experiments.
    $\bm{\varepsilon=1.0}$ $\bm{\varepsilon=0.3}$ $\bm{\varepsilon=0.1}$
    2dd, Sinkhorn-divergence 1.24 0.99 1.02
    $\bm{\varepsilon=}$1e-3 $\bm{\varepsilon=}$3e-4 $\bm{\varepsilon=}$1e-4
    Data Aggregation (Poisson), $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ 0.012 0.008 0.008
    Table 11: Effect of batch size in the 2dd and data aggregation experiments.
    Batch size 64 246 1024 4096
    2dd, Sinkhorn-divergence 1.57 1.24 1.04 1.04
    Data Aggregation (Poisson), $\mathrm{BW}_{2}^{2}\mathrm{-UVP}$ 0.032 0.017 0.013 0.012
    E.2 Ave! Celeba benchmark
    In this section, we provide an example that illustrates a potential limitation of our approach. As previously discussed, it is difficult to evaluate performance of barycentre algorithms in high dimensions due to the lack of a ground-truth. To combat this, [47] proposed the Ave, celeba! barycenter benchmark, which consists of 3 distributions of transformed CelebA faces [58], for which the $(\tfrac{1}{4},\tfrac{1}{2},\tfrac{1}{4})$-Wasserstein-2 barycentre recovers the true CelebA dataset. The resulting dataset consists of around 67k samples in each marginal, and each image is shape $64\times 64\times 3$.
    We consider applying the TreeDSBM in this example. It is known that performing bridge-matching between complex datasets such as images can be challenging, so for the first step we instead pretrain models using single bridge-matching iteration from a standard Gaussian to each marginal. To obtain the next coupling for training, we run the process from one of of the marginals to the latent representation in the Gaussian, and then out to the other marginals. This aids in learning, as there is often good structure preserved between the obtained samples from each marginal. For subsequent iterations we also warmstart the parameters from these pretrained models. The experiments were conducted on Nvidia A100 GPUs on Google Colaboratory.
    Hyperparameters
    We use the UNet architecture of [82], with 128 channels, channel multiples of (1,2,2,2), attention at layers (32,16,8), and 4 residual blocks at each layer.
    We use $\sigma=0.01$ and train each bridge-matching procedure with batch size 32, at a learning rate of 1e-4 and with exponential moving average weight of 0.999. For pretraining, we run for 20,000 training steps (which takes approximately 5 hours), and for subsequent IMF iterations we run for 10,000 steps (which each take around 2.5 hours).
    Discussion of results
    TreeDSBM is able to scale to the high-dimensional setting, but the obtained samples do not match the visual quality of state-of-the-art results such as those reported in [45]. Observe that some of the generated samples are good, but some contain additional artifacts that should not be present. We anticipate that this is due to the initial pretraining coupling. When generating samples $Y_{i}$ from the pretraining coupling, we simulate from one marginal to the Gaussian latent, and then out to the other marginals. This often results in strong structural similarities between the obtained $Y_{i}$ which yields good barycentre samples for training the next IMF iteration. However, sometimes the coupling samples $Y_{i}$ do not resemble each other, and the resulting barycentre sample consists of separate overlaid images. This appears to be difficult for the algorithm to recover from, resulting in the artifacts visible in some of the generated images.
    Overall this suggests a limitation of TreeDSBM for this particular benchmark—the true transport maps are in fact very simple in this specific example (primarily just colour changes), but it is difficult for TreeDSBM to learn this because communication between the edges is infrequent and only occurs after each IMF iteration. In contrast, state-of-the-art methods for this benchmark optimise all the maps together and with much more interaction between them, which we anticipate is a better inductive bias for the shared structure present in this benchmark. Note that the fact that TreeDSBM optimises the edges separately is in fact a strength of the approach in many settings; it results in stable training without needing adversarial objectives, and allows for speed-ups by training the edges simultaneously. However, this experiment suggests that this may be a limitation of our method in scenarios where the true maps exhibit a lot of shared structure (as in the case in this example), as communication between edges occurs too infrequently to recognise this shared structure.
    We remark that we observed improved performance in this benchmark by using a shared architecture over the edges (conditioning on the edge and the direction), compared to using a different network along each edge. Such an architecture makes sense in this example, as there are shared features along the edges that the network can learn and this reduces the computational and memory cost. When using a shared architecture, the network also appeared to create more consistent samples from the initial coupling. However, there are still unwanted artifacts present in many of the generated samples, and so alternative methods such as [45] are likely more suitable for settings such as these, as discussed above.
    We anticipate that the performance of TreeDSBM in this setting could be improved through architectural changes and other implementation tricks from the flow-matching literature (for example, using preconditioned flow parameterisations [41], and techniques to mitigate marginal drift in reflow methods [42]). Such investigations offer promising directions for future work.
    Figure 8: Samples from TreeDSBM applied to the Ave! Celeba benchmark. Many samples are transported well, but some pick up unwanted artifacts. We discuss the findings from this experiment below.
    E.3 Beyond star-shaped trees
    So far, we have demonstrated the empirical performance of TreeDSBM only on star-shaped trees, as we have focused on computing Wasserstein barycentres. Finally, we demonstrate that our TreeDSBM also works for non-star shaped trees, and thus has potential applications beyond only barycentre computation. We consider a simple 2-dimensional example with the same tree structure as shown in the TreeIMF diagram in Figure 4, with standard scikit-learn distributions on the observed leaves and each edge having length 1. As in the previous 2-dimensional barycentre problem, this is a challenging task due to the discontinuous transport maps and fragmented supports at the solution. We run TreeDSBM for 4 IMF iterations with $\varepsilon=0.1$, and train each edge for 10,000 iterations with learning rate 1e-3 and exponential moving average parameter of 0.99. We plot the obtained measures in Figure 9, and see that TreeDSBM is again able to learn the complex mappings required for this setting. While direct applications of non-star-shaped trees are less clear than in the barycentre case, examples have been studied in [38, 79], and they could have potential applications for modelling temporal behaviour of population dynamics, for example if populations were known to split according to a known structure. We leave investigating possible applications of general tree-structured costs for future work.
@@ Appendix F Licenses
    NeurIPS Paper Checklist
    1. 1.
      Claims
    2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
    3. Answer: [Yes]
    4. Justification: In the abstract and introduction, we state that we extend the IMF procedure to the tree-based SB setting. In Section 3 we demonstrate the theoretical soundness of our proposed approach, and we explain how to implement it in Section 3.3. In Section 4 we demonstrate the empirical performance of our approach, showing it inherits the benefits on IMF over IPF in the tree-based setting.
    5. 2.
      Limitations
    6. Question: Does the paper discuss the limitations of the work performed by the authors?
    7. Answer: [Yes]
    8. Justification: We discuss the limitations of our approach in Section 5. We also provide an experiment illustrating a potential limitation of our approach in Appendix E.
    9. 3.
      Theory assumptions and proofs
    10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
    11. Answer: [Yes]
    12. Justification: We present simple statements of the theoretical results in the main text, with full assumptions and proofs included in Appendix B.
    13. 4.
      Experimental result reproducibility
    14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
    15. Answer: [Yes]
    16. Justification: We report full implementation details of our reported experiments in Appendix C for reproducibility, including neural architecture and hyperparameter choices and data generation procedures.
    17. 5.
      Open access to data and code
    18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
    19. Answer: [Yes]
    20. Justification: We provide code to run our experiments in the supplementary material. For alternative algorithms, we use the open-source code provided by authors to which we provide links in Appendix F.
    21. 6.
      Experimental setting/details
    22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
    23. Answer: [Yes]
    24. Justification: We include these details in Appendix D.
    25. 7.
      Experiment statistical significance
    26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
    27. Answer: [No]
[…]
    29. 8.
      Experiments compute resources
    30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
    31. Answer: [Yes]
    32. Justification: These details are included in Appendix D.
    33. 9.
      Code of ethics
    34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
    35. Answer: [Yes]
    36. Justification: We have reviewed the Code of Ethics and confirm that our work conforms to these guidelines. We see no potential harmful consequences of our work, and include extensive reproducibility details in Appendix D.
    37. 10.
      Broader impacts
    38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
    39. Answer: [N/A]
    40. Justification: Our work is mostly theory and methodology focused for the tree-structured Schödinger Bridge problem. We do not see any immediate societal impacts, though certain applications may share similar societal consequences as in flow-based generative models upon which our approach is based.
    41. 11.
      Safeguards
    42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
    43. Answer: [N/A]
    44. Justification:
    45. Guidelines: Our work poses no such risks.
    46. 12.
      Licenses for existing assets
    47. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
    48. Answer: [Yes]
    49. Justification: We cite the works that we use, and include the licenses of code and datasets in Appendix F.
    50. 13.
      New assets
    51. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
    52. Answer: [N/A]
    53. Justification: The paper does not release new assets.
    54. 14.
      Crowdsourcing and research with human subjects
    55. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
    56. Answer: [N/A]
    57. Justification: The paper does not involve crowdsourcing nor research with human subjects.
    58. 15.
      Institutional review board (IRB) approvals or equivalent for research with human subjects
    59. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
    60. Answer: [N/A]
    61. Justification: The paper does not involve crowdsourcing nor research with human subjects.
    62. 16.
      Declaration of LLM usage
    63. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
    64. Answer: [N/A]
    65. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.